OpenStudy (anonymous):

How would I graph f(x)=27(1.3)x? (It's a exponential function so it has to be a curved line?)

OpenStudy (anonymous):

I made a chart to help me understand but I'm not sure how.. x|f(x) ----- 0|27 1|- 2|- 3|- ---

OpenStudy (anonymous):

Would I just do like 27*1.3*2?

OpenStudy (afrodiddle):

well let's fill in the chart

OpenStudy (anonymous):

would 1=35.1?

OpenStudy (afrodiddle):

27(1.3) ^1

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so it would be 35.1

OpenStudy (afrodiddle):

Yes

OpenStudy (anonymous):

First question...Is itf(x) = 27(1.3)^x or f(x)=27(1.3)x? f(x)=27(1.3)x f(1) = 27(1.3)(1) f(2) = 27(1.3)(2)

OpenStudy (anonymous):

then 2 would be 45.63?

OpenStudy (afrodiddle):

It has to be 27(1.3) ^ x Other wise it's not exponential.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Oh very true!

OpenStudy (anonymous):

would 2 be 45.63?

OpenStudy (afrodiddle):

yes

OpenStudy (anonymous):

then 3 would be 59.319?

OpenStudy (anonymous):

so then i would plot like (0,27), (1,35.1) (2,45.63) and (3,59.319) with a curved line right?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

thank you :)